Circular No. 7421

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cf IN MCG +11-19-25 T. Puckett, Mountain Town, GA; and A. Sehgal, Woodinville, WA,

report their discovery of an apparent supernova (mag 16.9) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory automated 0.60-m supernova patrol telescope on May 9.23 UT. SN 2000cf is located at α = 15^{h}52^{m}56^{s}.19, δ = +65^{o}56′12′′.7 (equinox 2000.0), which is 3′′.1 east and 4′′.3 north of the center of MCG +11-19-25. The new object was also present on an unfiltered CCD frame taken on May 10.05, but it was not present on a frame taken on Mar. 7.36, and it does not appear on Palomar Sky Survey images taken on 1994 June 1 (limiting mag ∼ 21.0) or 1955 May 17 (limiting mag ∼ 20.0).

SUPERNOVA 2000cg IN UGC 10121 Y. Sato and W. D. Li, University of California at Berkeley, on behalf

of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope of an apparent supernova on unfiltered images taken on May 6.4 (mag ∼ 18.6) and 9.4 UT (mag ∼ 19.0, limiting mag ∼ 20.5). The new object also appeared on an image taken on Apr. 27.4 UT (mag ∼ 18.6). SN 2000cg is located at α = 15^{h}59^{m}45^{s}.52, δ = +18^{o}48′13′′.0 (equinox 2000.0), which is 4′′.0 west and 9′′.0 north of the nucleus of UGC 10121. Unfiltered images of the field taken on 2000 Apr. 1.4 (limiting mag ∼ 19.0) and 1998 July 21.3 (limiting mag ∼ 20.0) showed nothing at this position.

SUPERNOVA 2000ch IN NGC 3432 A. V. Filippenko, University of California at Berkeley, writes: “The

descriptions of the ‘variable star’ in NGC 3432 (cf. IAUC 7415, 7417, 7419) strongly suggest that it is a subluminous type-IIn supernova similar to SN 1997bs in NGC 3627 (cf. IAUC 6627) and SN 1999bw in NGC 3198 (cf. IAUC 7152). The progenitors of these subluminous objects may be very massive, powerful stars such as luminous blue variables (e.g., SN 1961V, cf. Goodrich et al. 1989, Ap.J. 342, 908; SN 1997bs, cf. Van Dyk et al. 1999, A.J. 118, 2331). If so, the ‘explosion mechanism’ might not be core collapse, and the star may survive (e.g., SN 1961V, cf. Filippenko et al. 1995, A.J. 110, 2261). Indeed, the ‘explosion’ is actually just a superoutburst of the luminous variable. Some of these outbursts last longer than others, and that of the object in NGC 3432 seems to have been short-lived.”

2000 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green